# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ 5.5   The role of externalisation and stressing model parameters
Figure 12 illustrates that, with the original filtering framework, the broker’s externalisation rate is centred around zero. In contrast, the median externalisation in the naive case (right panel) is significantly higher, ranging between $70$ and $90\%$. This discrepancy is due to the difference in signal estimation quality between $\hat{\alpha}$ and $\hat{\alpha}^{\text{naive}}$. The ordering of the lines when stressing $\kappa^{\alpha}$ in the right panel of Figure 12 is similar to those in Figure 10.
Appendix A Proof of Proposition 1
@@
$\begin{split}\mathrm{d}Z_{t}&=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\end{split}$ (A.1)
@@
$\mathrm{d}\tilde{W}_{t}:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]$ (A.2)
@@
$I_{t}:=Z_{t}-\int_{0}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(s)\,\alpha_{s}\,\Big{|}\,\mathcal{Z}_{s}\right]\,\mathrm{d}s$ (A.3)
@@
$\begin{split}\mathbb{E}\left[I_{t}-I_{s}\,|\,\mathcal{Z}_{s}\right]&=\mathbb{E}\left[Z_{t}-Z_{s}-\int_{s}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(u)\,\alpha_{u}\,\Big{|}\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\Big{|}\,\mathcal{Z}_{s}\right]\\ &=\mathbb{E}\left[Z_{t}-Z_{s}\,|\,\mathcal{Z}_{s}\right]-\mathbb{E}\left[\int_{s}^{t}\mathfrak{G}_{7}(u)\,\alpha_{u}\,\mathrm{d}u\,\Big{|}\,\mathcal{Z}_{s}\right]\,.\end{split}$ (A.4)
Note that the diffusion terms in (A.2) are bounded by $1$, so that
@@
But $\mathcal{Z}_{s}\subset\mathcal{F}_{s}$, so that $\mathbb{E}[Z_{t}-Z_{s}\,|\mathcal{Z}_{s}]=\mathbb{E}[\mathbb{E}[Z_{t}-Z_{s}\,|\,\mathcal{F}_{s}]\,|\,\mathcal{Z}_{s}]$, and we obtain (A.4) equals zero.
@@
$M_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,,$ (A.5)
$N_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u+\int_{0}^{t}\sigma^{\alpha}\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,$ (A.6)
@@
$\mathbb{E}\left[M_{t}-M_{s}\,\big{|}\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]=0\,.$ (A.7)
@@
$\mathbb{E}\left[N_{t}-N_{s}\,\big{|}\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]=0\,.$ (A.8)
@@
$\hat{\alpha}^{\text{alt}}_{t}=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]=M_{t}-\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u=\int_{0}^{t}r_{u}^{M}\,\mathrm{d}I_{u}\ -\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u\,.$ (A.9)
@@
$\mathrm{d}(\hat{\alpha}^{\text{alt}}_{t})^{2}=\left[-2\,\kappa^{\alpha}\,(\hat{\alpha}^{\text{alt}}_{t})^{2}+(r_{t}^{M})^{2}\right]\,\mathrm{d}t+2\,r_{t}^{M}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}I_{t}\,.$ (A.10)
@@
$\begin{split}\mathbb{E}\left[\alpha_{t}^{2}\,\big{|}\,\mathcal{Z}_{t}\right]&=\int_{0}^{t}\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big{|}\,\mathcal{Z}_{t}\right]\Big{\}}\,\mathrm{d}u+\sigma^{\alpha}\,\mathbb{E}\left[\int_{0}^{t}2\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]\\ &=\int_{0}^{t}2\,r_{u}^{N}\,\mathrm{d}I_{u}+\int_{0}^{t}\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big{|}\,\mathcal{Z}_{u}\right]\Big{\}}\,\mathrm{d}u\,.\end{split}$ (A.11)

### v2

@@ A.1  Hamilton–Jacobi–Bellman equation
to obtain that $h^{I}$ satisfies the partial differential equation (PDE)
with terminal condition
Next, use the ansatz
to obtain the system of differential equations
$0$ $=-(\rho_{0}+\rho_{1}\mathbb{V}^{I}_{t})+(g_{2}^{I})^{\prime}+\frac{(g_{2}^{I})^{2}}{\mathfrak{b}}$ (A.4)
$0$ $=\partial_{t}\,g_{1}^{I}+(\mathfrak{p}\hat{\nu}+\alpha)+\frac{g_{1}^{I}\,g_{2}^{I}}{2\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{1}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{1}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{1}^{I}\,$,
$\phantom{{}={}}+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\,\hat{\nu}}\,g_{1}^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\rho\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\partial_{\alpha\,\hat{\nu}}\,g_{1}^{I}\,$, (A.5)
$0$ $=\partial_{t}\,g_{0}^{I}+\frac{(g_{1}^{I})^{2}}{4\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{0}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{0}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{0}^{I}$
$\phantom{{}={}}+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\,\hat{\nu}}\,g_{0}^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\rho\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\partial_{\alpha\,\hat{\nu}}\,g_{0}^{I}\,$, (A.6)
with terminal conditions
$g_{2}^{I}(T)\quad$ $=-(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{T})\,$, (A.7)
[…]
The following lemma solves the system of PDEs (A.4)-(A.5)-(A.6).
Lemma 1.
[…]
solve the system of differential equations (A.6) with terminal conditions (A.7)-(A.8)-(A.9).
Proof.
Observe that (A.4) is a one-dimensional Riccati equation with quadratic term ${1}/{\mathfrak{b}}>0$, “constant” term $(\rho_{0}^{I}+\rho_{1}^{I}\,\mathbb{V}^{I}_{t})<0$, and terminal value $(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{t})<0$. Use Lemma 4.1 in Reid (1972) to obtain that the equation has a unique solution in $(-\infty,T]$, and that it is negative.
Next, to solve (A.5) and (A.6), one may work backwards by substituting the desired form of $g_{0}$ and $g_{1}$ into the system of equations to obtain
$0$ $=\alpha$ $\left[1-\kappa^{\alpha}\,z_{1}^{I}+\frac{g_{2}^{I}\,z_{1}^{I}}{2\,\mathfrak{b}}+(z_{1}^{I})^{\prime}\right]$ (A.12)
$\phantom{{}={}}+\hat{\nu}$ $\left[\mathfrak{p}-\theta^{B}z_{2}^{I}+\frac{g_{2}^{I}\,z_{2}^{I}}{2\,\mathfrak{b}}+(z_{2}^{I})^{\prime}\right]\,$, (A.13)
$0$ $=$ $\left[\frac{\mathfrak{p}\,\sigma^{\alpha}\,\rho\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\,z_{6}^{I}+(\sigma^{\alpha})^{2}\,z_{7}^{I}+\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}z_{8}^{I}+(z_{3}^{I})^{\prime}\right]$
$\phantom{{}={}}+\hat{\nu}$ $\left[\frac{g_{2}^{I}\,z_{2}^{I}}{2\,\mathfrak{b}}-\theta^{B}\,z_{5}^{I}+(z_{5}^{I})^{\prime}\right]+\alpha\left[\frac{g_{2}^{I}\,z_{1}^{I}}{2\,\mathfrak{b}}-\kappa^{\alpha}z_{4}^{I}+(z_{4}^{I})^{\prime}\right]$
$\phantom{{}={}}+\alpha\,\hat{\nu}$ $\left[\frac{z_{1}^{I}\,z_{2}^{I}}{2\,\mathfrak{b}}-\kappa^{\alpha}\,z_{6}^{I}-\theta^{B}\,z_{6}^{I}+(z_{6}^{I})^{\prime}\right]+\alpha^{2}\left[\frac{(z_{1}^{I})^{2}}{4\,\mathfrak{b}}-2\,\kappa^{\alpha}\,z_{7}^{I}+(z_{7}^{I})^{\prime}\right]+\hat{\nu}^{2}\left[\frac{(z_{2}^{I})^{2}}{4\,\mathfrak{b}}-2\,\theta^{B}\,z_{8}^{I}+(z_{8}^{I})^{\prime}\right]\,.$ (A.14)
$(z_{1}^{I},z_{2}^{I})\to(z_{6}^{I},z_{7}^{I},z_{8}^{I})\to(z_{4}^{I},z_{5}^{I})\to z_{3}^{I}\,.$ (A.15)
In the above order, the ODEs are linear, so there exist $\{z_{i}^{I}\}_{i=1,\cdots,8}$ satisfying (A.5) and (A.6). ∎
A.2  Proof of Theorem 1
Proof.
Observe that (3.9) satisfies the HJB equation in (A.2). Moreover, by substituting $g_{i}^{I}$ for $i=0,1,2$ from Lemma 1, the solution to the HJB equation is a quadratic polynomial in terms of $\mathfrak{x}$ with time-dependent $C^{1}([0,T])$ coefficients; thus, the solution is bounded by a quadratic growth. By Theorem 3.5.2 (Verification Theorem) in Pham (2009), the trader’s strategy (3.11) is indeed the optimal Markovian control. The last part of the theorem is as follows: $g_{2}^{I}$ is negative from Lemma 1.
Next, by solving for the ODE (A.13), we obtain
$z_{2}^{I}(t)=\int_{t}^{T}\mathfrak{p}\,\exp\left\{\int_{t}^{s}\frac{g_{2}^{I}(u)}{2\,\mathfrak{b}}-\theta^{B}\,\mathrm{d}u\right\}\,\mathrm{d}s\geq 0\quad\forall t\in[0,T]\,$, (A.16)
which shows that $z_{2}^{I}$ is non-negative. The same argument follows for $z_{1}^{I}$ by solving (A.12). Observe that the equation above is linear in $\mathfrak{p}$. ∎
Appendix B Derivations for the control problem of the broker
B.1  Filtering the order flow
We compute the dynamics of ${\gamma}_{t}$ to obtain
$\begin{split}\mathrm{d}{\gamma}_{t}&=\quad\left[f_{1}^{\prime}(t)\,\alpha_{t}-\kappa^{\alpha}\,f_{1}(t)\,\alpha_{t}+f_{2}^{\prime}(t)\,\hat{\nu}_{t}-\theta^{B}\,f_{2}(t)\,\hat{\nu}_{t}\right]\,\mathrm{d}t+\left[\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{f_{2}(t)}{(\sigma^{S})^{2}}\,\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\left(\mathrm{d}Y_{t}-\mathfrak{p}\,\hat{\nu}_{t}\,\mathrm{d}t\right)\right]\\ &=\quad\Bigg[\Big(f_{1}^{\prime}(t)-\kappa^{\alpha}\,f_{1}(t)\Big)\,\alpha_{t}+\left(f_{2}^{\prime}(t)-\theta^{B}\,f_{2}(t)-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\right)\,\hat{\nu}_{t}+\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\,\nu_{t}\Bigg]\,\mathrm{d}t\\ &\hskip 227.62204pt+\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,\mathrm{d}W_{t}^{S}.\end{split}$ (B.1)
The broker knows that
so it follows that
[…]
for functions $\mathfrak{G}_{i}$ with $i\in\{0,1,2,3,4\}$ and $\mathfrak{G}^{\alpha}$ given by
$\mathfrak{G}_{0}(t)$ $=-\theta^{B}-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}+\frac{f_{2}^{\prime}(t)}{f_{2}(t)}$ (B.3)
$=-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}-\frac{\mathfrak{p}}{2\,f_{2}(t)\,\mathfrak{b}}-\frac{f_{3}(t)}{2}\,$, (B.4)
$\mathfrak{G}^{\alpha}(t)$ $=f_{1}^{\prime}(t)-f_{1}(t)\,\left(\kappa^{\alpha}+\mathfrak{G}_{0}(t)\right)$ (B.5)
$=-\frac{1}{2\,\mathfrak{b}}-f_{1}(t)\,\left(\frac{f_{3}(t)}{2}+\mathfrak{G}_{0}(t)\right)\,$, (B.6)
$\mathfrak{G}_{1}(t)$ $=\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}},\quad\mathfrak{G}_{2}(t)=0\,$, (B.7)
$\mathfrak{G}_{3}(t)$ $=\sigma^{\alpha}\,f_{1}(t),\quad\quad\mathfrak{G}_{4}(t)=\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,.$ (B.8)
We define the process $\tilde{Z}_{t}:=\gamma_{t}/\mathfrak{G}_{5}(t)$ with $\mathfrak{G}_{5}(t):=\sqrt{\mathfrak{G}_{3}(t)^{2}+\mathfrak{G}_{4}(t)^{2}+2\,\rho\,\mathfrak{G}_{3}(t)\,\mathfrak{G}_{4}(t)}$ and obtain
$\begin{split}\mathrm{d}\tilde{Z}_{t}&=\quad\left[\frac{-\tilde{Z}_{t}\,\mathfrak{G}_{5}^{{}^{\prime}}(t)}{\mathfrak{G}_{5}(t)}+\frac{\mathfrak{G}^{\alpha}(t)}{\mathfrak{G}_{5}(t)}\,\alpha_{t}+\frac{\mathfrak{G}_{0}(t)}{\mathfrak{G}_{5}(t)}\,\gamma_{t}+\frac{\mathfrak{G}_{1}(t)}{\mathfrak{G}_{5}(t)}\,\nu_{t}+\frac{\mathfrak{G}_{2}(t)}{\mathfrak{G}_{5}(t)}\right]\,\mathrm{d}t+\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,.\end{split}$ (B.9)
The coefficients in the above expression are well-defined when $t\to T$ and we have that
[…]
where from L’Hopital’s rule and the differential equations (A.12)–(A.13), it follows that $\lim_{t\uparrow T}{f_{1}(t)}/{\mathfrak{G}_{5}(t)}$ and $\lim_{t\uparrow T}{f_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite. This implies that $\tilde{Z}$ is a continuous square-integrable process. Similarly, one also shows that the limits of ${\mathfrak{G}^{\alpha}(t)}/{\mathfrak{G}_{5}(t)}$, ${\mathfrak{G}_{1}(t)}/{\mathfrak{G}_{5}(t)}$, and ${\mathfrak{G}_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite as $t\uparrow T$. On the other hand, observe that ${\mathfrak{G}_{5}^{{}^{\prime}}(t)}/{\mathfrak{G}_{5}(t)}$ and ${\mathfrak{G}_{0}(t)}/{\mathfrak{G}_{5}(t)}$ blow up as $t\uparrow T$. However, because $(\tilde{Z}_{t})_{t\in{\mathfrak{T}}}$ and $\big(\int_{0}^{t}\alpha_{s}\,{\mathfrak{G}^{\alpha}(s)}/{\mathfrak{G}_{5}(s)}+{\mathfrak{G}_{1}(s)}/{\mathfrak{G}_{5}(s)}\,\nu_{s}+{\mathfrak{G}_{2}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big)_{t\in{\mathfrak{T}}}$ are linear combinations of square-integrable progressively measurable processes, the process $\big(\int_{0}^{t}{-\tilde{Z}_{s}\,\mathfrak{G}_{5}^{{}^{\prime}}(s)}/{\mathfrak{G}_{5}(s)}+\gamma_{s}\,{\mathfrak{G}_{0}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big)_{t\in{\mathfrak{T}}}$ is also square-integrable and progressively measurable. Next we define the main process that we use in the filtering framework of this section. The broker observes the process $Z_{t}$ defined as
[…]
$\mathrm{d}\,\tilde{W}_{t}$ $:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,$, (B.12)
$\mathfrak{G}_{6}(t)$ $=-\mathfrak{G}_{5}^{{}^{\prime}}(t)/\mathfrak{G}_{5}(t)\,$, (B.13)
$\mathfrak{G}_{7}(t)$ $=\mathfrak{G}^{\alpha}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{8}(t)=\mathfrak{G}_{0}(t)/\mathfrak{G}_{5}(t)\,$, (B.14)
$\mathfrak{G}_{9}(t)$ $=\mathfrak{G}_{1}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{10}(t)=\mathfrak{G}_{2}(t)/\mathfrak{G}_{5}(t)\,.$ (B.15)
Use the ODEs (A.12), (A.13), and (3.4), to rewrite $\mathfrak{G}_{6}(t)$ as
$\mathfrak{G}_{6}(t)=-\frac{\mathfrak{G}_{3}^{{}^{\prime}}(t)[\mathfrak{G}_{3}(t)+\rho\,\mathfrak{G}_{4}(t)]+\mathfrak{G}_{4}^{{}^{\prime}}(t)[\rho\,\mathfrak{G}_{3}(t)+\mathfrak{G}_{4}(t)]}{\mathfrak{G}_{5}(t)^{2}}\,$, (B.16)
where
$\mathfrak{G}_{3}^{{}^{\prime}}(t)$ $=\sigma^{\alpha}\left[-\frac{1}{2\,\mathfrak{b}}+\kappa^{\alpha}\,f_{1}(t)-\frac{f_{3}(t)\,f_{1}(t)}{2}\right]\,$, (B.17)
$\mathfrak{G}_{4}^{{}^{\prime}}(t)$ $=\frac{\mathfrak{p}}{\sigma^{S}}\left[\mathbb{V}^{I}_{t}\,\left(-\frac{\mathfrak{p}}{2\,\mathfrak{b}}+\theta^{B}\,f_{2}(t)-\frac{f_{3}(t)\,f_{2}(t)}{2}\right)+f_{2}(t)\,\left((\sigma^{B})^{2}-2\,\theta^{B}\,\mathbb{V}^{I}_{t}-\frac{\mathfrak{p}^{2}\,(\mathbb{V}^{I}_{t})^{2}}{(\sigma^{S})^{2}}\right)\right]\,.$ (B.18)
B.2  Proof of Proposition 1
@@
$\begin{split}\mathrm{d}Z_{t}&=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\end{split}$ (B.19)
@@
$\mathrm{d}\tilde{W}_{t}:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]$ (B.20)
@@
$I_{t}:=Z_{t}-\int_{0}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(s)\,\alpha_{s}\,\Big|\,\mathcal{Z}_{s}\right]\,\mathrm{d}s$ (B.21)
@@
$\begin{split}\mathbb{E}\left[I_{t}-I_{s}\,|\,\mathcal{Z}_{s}\right]&=\mathbb{E}\left[Z_{t}-Z_{s}-\int_{s}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(u)\,\alpha_{u}\,\Big|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\Big|\,\mathcal{Z}_{s}\right]\\ &=\mathbb{E}\left[Z_{t}-Z_{s}\,|\,\mathcal{Z}_{s}\right]-\mathbb{E}\left[\int_{s}^{t}\mathfrak{G}_{7}(u)\,\alpha_{u}\,\mathrm{d}u\,\Big|\,\mathcal{Z}_{s}\right]\,.\end{split}$ (B.22)
Note that the diffusion terms in (B.20) are bounded by $1$, so that
@@
But $\mathcal{Z}_{s}\subset\mathcal{F}_{s}$, so that $\mathbb{E}[Z_{t}-Z_{s}\,|\mathcal{Z}_{s}]=\mathbb{E}[\mathbb{E}[Z_{t}-Z_{s}\,|\,\mathcal{F}_{s}]\,|\,\mathcal{Z}_{s}]$, and we obtain (B.22) equals zero.
@@
$M_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big|\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,$, (B.23)
$N_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u+\int_{0}^{t}\sigma^{\alpha}\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big|\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,$ (B.24)
@@
$\mathbb{E}\left[M_{t}-M_{s}\,\big|\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u\,\big|\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big|\,\mathcal{Z}_{s}\right]=0\,.$ (B.25)
@@
$\mathbb{E}\left[N_{t}-N_{s}\,\big|\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u\,\big|\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big|\,\mathcal{Z}_{s}\right]=0\,.$ (B.26)
@@
$\hat{\alpha}^{\text{alt}}_{t}=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big|\,\mathcal{Z}_{t}\right]=M_{t}-\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u=\int_{0}^{t}r_{u}^{M}\,\mathrm{d}I_{u}\ -\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u\,.$ (B.27)
@@
$\mathrm{d}(\hat{\alpha}^{\text{alt}}_{t})^{2}=\left[-2\,\kappa^{\alpha}\,(\hat{\alpha}^{\text{alt}}_{t})^{2}+(r_{t}^{M})^{2}\right]\,\mathrm{d}t+2\,r_{t}^{M}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}I_{t}\,.$ (B.28)
@@
$\begin{split}\mathbb{E}\left[\alpha_{t}^{2}\,\big|\,\mathcal{Z}_{t}\right]&=\int_{0}^{t}\Big\{(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big|\,\mathcal{Z}_{t}\right]\Big\}\,\mathrm{d}u+\sigma^{\alpha}\,\mathbb{E}\left[\int_{0}^{t}2\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big|\,\mathcal{Z}_{t}\right]\\ &=\int_{0}^{t}2\,r_{u}^{N}\,\mathrm{d}I_{u}+\int_{0}^{t}\Big\{(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big|\,\mathcal{Z}_{u}\right]\Big\}\,\mathrm{d}u\,.\end{split}$ (B.29)